### v1

# What could go wrong? Discovering and describing failure modes in computer vision

Abstract
Deep learning models are effective, yet brittle. Even carefully trained, their behavior tends to be hard to predict when confronted with out-of-distribution samples. In this work, our goal is to propose a simple yet effective solution to predict and describe via natural language potential failure modes of computer vision models. Given a pretrained model and a set of samples, our aim is to find sentences that accurately describe the visual conditions in which the model underperforms. In order to study this important topic and foster future research on it, we formalize the problem of Language-Based Error Explainability (LBEE) and propose a set of metrics to evaluate and compare different methods for this task. We propose solutions that operate in a joint vision-and-language embedding space, and can characterize through language descriptions model failures caused, e.g., by objects unseen during training or adverse visual conditions. We experiment with different tasks, such as classification under the presence of dataset bias and semantic segmentation in unseen environments, and show that the proposed methodology isolates nontrivial sentences associated with specific error causes. We hope our work will help practitioners better understand the behavior of models, increasing their overall safety and interpretability.
1 Introduction
The sharp contrast between the ideal conditions found in standard benchmarks and the unpredictable nature of the real world majorly hinders the deployment of computer vision systems in the wild. Despite the most meticulous efforts, samples used to train and validate visual models will only represent a fraction of the diversity that these models will face once deployed. It is thus critical to detect model vulnerabilities and bring them to the user’s attention in an interpretable way [6].
Among research lines that focus on advancing model interpretability, a prominent one is that of explaining errors made by computer vision models via natural language descriptions. Different works have addressed the problem of Language-Based Error Explainability (LBEE) before [9, 51, 20, 49, 34]. Yet, none of them proposed ways to quantitatively assess the predicted descriptions, mainly confining them to a qualitative inspection of the model’s error modes. We assert that the LBEE problem requires appropriate metrics, in order to rank methods and track progress in this field. In this paper, we take different steps in this direction through the following contributions: i) we provide a rigorous formalization of the LBEE problem, ii) we propose a family of methods to tackle it, and iii) we introduce different metrics to evaluate performance.
In short, the problem at hand is the following: Given a set of images, denoted by $\mathcal{X}$, a task-specific pre-trained model $\mathcal{M}_{\theta}$, and a large sentence set ${\mathcal{S}}$, our desired output is a subset of sentences ${\mathcal{S}}^{*}\subset{\mathcal{S}}$ that describes in plain text the samples from $\mathcal{X}$ on which the model $\mathcal{M}_{\theta}$ under-performs. To tackle this problem, we propose a simple, yet powerful family of approaches that summarize the visual conditions deemed challenging for a given model via natural language. Our methods rely on uncertainty estimation to distinguish between samples for which the model is more likely to perform well or poorly, and on clustering in a joint vision-and-language embedding space to group together challenging samples associated with specific visual conditions. Operating in this joint representation space allows us to associate sentences to each identified cluster of samples (see illustration in Fig. 1). In order to isolate sentences that are unique to each cluster composed of under-performing samples, we propose different ways to contrast descriptions of those clusters on which the model struggles with those on which it performs performs well (denoted as hard and easy clusters, respectively).
Figure 1: Overview. Provided with a pretrained model ${\mathcal{M}}_{\theta}$, a target image set ${\mathcal{X}}$, and a set of sentences ${\mathcal{S}}$, the family of solutions we propose for LBEE follows the following steps. In Step 1, images from the target set ${\mathcal{X}}$ are split into easy and hard samples, based on the model’s confidence. In Step 2, samples are embedded using Open-CLIP and the visual embeddings of each set are clustered independently. In Step 3, for each hard and easy cluster we compute the cosine similarities between the textual embeddings of candidate sentences and the cluster prototypes. We also assign to each hard prototype the closest easy prototype. Step 4 performs sentence selection for that hard cluster, based on these sentence-prototype similarities. Step 5 aggregates the cluster-specific sentence sets to produce the output.
Our proposed family of approaches is unsupervised and task-agnostic, departing from previous methods which have been specifically designed for classification [9, 51, 20], and often require access to ground-truth or privileged information, such as the types of errors to be discovered [9, 51]. To assess how well these methods find suitable failure explanations, we introduce metrics specifically designed for LBEE capturing whether or not the sentences satisfactorily explain reasons for failure, and how well the produced sentences describe the images within each error mode.
While the need to provide the sentence set ${\mathcal{S}}$ is a practical limitation of our methods, it plays a crucial role as it allows 1) defining a more generic ground-truth (GT) for evaluation that goes beyond human-in-the-loop qualitative assessment, 2) finding more subtle explanations than clear test set’s biases (e.g.“waterbirds in land” [26], “blonde males” in CelebA [28] or human-implanted biases in NICO++{++}+ +  [9, 51]) 3) performing a fair and quantitative comparisons of different methods, 4) avoiding the need to involve the assessment with a large language model (as e.g.in  [8]).
In summary, the contributions of our paper are: i) formalizing the LBEE problem (Sec. 3); ii) proposing a family of methods to tackle it (Sec. 4); iii) introducing different metrics to evaluate performance; (Sec. 5) iv) showing the effectiveness of our proposed solutions through extensive experiments, focusing on both classification and segmentation segmentation tasks (Sec. 6).
2 Related work
Automated error discovery. Recently, methods to assess errors made by machine learning models have shifted from human-in-the-loop approaches [50, 42, 19, 5, 10] to automatic detection of consistent model failure modes [9, 51, 20, 49, 34].
In particular, Eyuboglu et al. [9, 51] propose to learn a mixture model to partition the data according to classification errors related to spurious correlations. Jain et al. [20] characterize model failures as directions in a latent space found with SVMs. Kim et al. [23] identify biases by analyzing captions from misclassified images with a Vision-Language Model (VLM) and ranks the keywords from the captions using CLIP [32]. Wiles et al. [49] use conditional text-to-image diffusion models to generate synthetic images that are grouped based on how the model misclassifies them. Rezaei et al. [34] extract human-understandable concepts (tags) with VLMs and examine the model’s behavior conditioning on the presence or absence of the combination of those tags. Also Metzen et al. [30] rely on text-to-image models: by encoding specific information in the prompt, they synthesize images from the combinatorially large set of data sub-populations, and run tests to find out on which ones the model under-performs. Finally, a recent work proposes methods to describe differences between two image sets [8]. While this method could be applied to LBEE, the proposed evaluation based on GPT-4 can only assess whether the difference between the sets is correct but not whether this difference is related or not to model failure.
In contrast with the above works, our method only require encoding the images and the textual explanations in a shared visual-textual space such as CLIP, where we rely on simple arithmetic operations among clusters to characterize them via natural language. Our solutions can be seamlessly applied to arbitrary image task—porting this line of work to semantic segmentation for the first time. Differently from previous work on this topic, we also provide a rigorous formalization of the LBEE problem, with ad hoc metrics.
Explainability. There is a growing body of work connecting explainable AI and CV. Several focus on producing visual explanations visualizing image regions that contribute most to model failures [35, 53, 40, 44]. A few works aim at generating counterfactual explanations [11, 16, 18], identifying small image edits that cause the model to fail. More recently, diffusion models have been used to this end [21, 2, 46]. In contrast to these methods which only explain failures on a particular image, our aim is to provide explanations of the model’s overall behavior on an entire image set, in order to uncover consistent error patterns and major weaknesses under particular conditions. Furthermore, we rely on natural language as a more interpretable interface.
Another body of work tackles interpretability by providing short text descriptions for the reasons a model has provided some output [13, 22, 15, 38]. Such methods are related to our work, in that we are also interested in generating language descriptions for model behaviors; yet, we are interested in failure modes, not in the general rationale behind a prediction.
3 Language-Based Error Explainability
We can define the LBEE problem as automatically finding and describing potential errors from a model based on its performance on a given dataset. Formally, let ${\mathcal{M}}_{\theta}$ be a computer vision model parameterized by $\theta$, for an arbitrary computer vision task, such as classification or segmentation. Let ${\mathcal{X}}=\{{\bf x}_{i}\}^{M}_{i=1}$ be a set of images sampled from an arbitrary target distribution, and let $\omega^{\text{avg}}_{\theta}$ be the average performance of the model on such data. Finally, let ${\mathcal{S}}=\{s_{n}\}^{N}_{n=1}$ be a broad and generic set of sentences that describe various visual elements and conditions, which can be either manually defined or generated with a large language model (LLM)—see Appendix C.
Problem formulation. Given a target set ${\mathcal{X}}$, our goal is to find a set of sentences ${\mathcal{R}}_{{\mathcal{S}}}\subset{\mathcal{S}}$ describing any likely failure causes for the model ${\mathcal{M}}_{\theta}$ with respect to images from ${\mathcal{X}}$—such as unseen objects, visual conditions never encountered during training, or errors of any other nature. Intuitively, each selected sentence in ${\mathcal{R}}_{{\mathcal{S}}}$ should describe elements present in images on which the models fails and not in those in which the model succeeds.
Letting $\omega^{s_{n}}_{\theta}$ be the average performance of the model on images for which the sentence $s_{n}$ is relevant, we determine that such sentence is describing a failure mode if the difference between the global average $\omega^{\text{avg}}_{\theta}$ (average over all images in ${\mathcal{X}}$) and $\omega^{s_{n}}_{\theta}$ is larger than a threshold $\beta$, with $\beta$ being a predefined margin. The desired output ${\mathcal{S}}^{*}_{\beta}$—used to evaluate the performance of the methods— is the collection of all of these sentences
${\mathcal{S}}^{*}_{\beta}=\{s_{n}\in{\mathcal{S}}|\,\omega_{\theta}^{s_{n}}<\omega_{\theta}^{\text{avg}}-\beta\}\enspace,$ (1)
Note that $\beta$ allows the user to set an arbitrary severity level for the error descriptions, i.e., a higher $\beta$ restricts the desired set ${\mathcal{S}}^{*}_{\beta}$ to more challenging sentences. We provide more details in Sec. C.2.
Relation with prior art. To conclude this section, we position the formulation introduced above with the one tackled by related works [9, 51, 20].
One core difference is that these methods limit their problem formulation to the classification task, while we do not make any assumption related to the nature of the task at hand. Furthermore, in [20] they analyze errors per class, and in [9, 51] assume access to ground truth or prior knowledge about the error types to look for (e.g., bias-conflicting sub-populations or context). In contrast to the above works, we do not use any knowledge about errors prior to evaluation and neither assume—in general—access to class information. Finally, while in prior works the language-based descriptions are part of the proposed methodology, they are not part of the problem formulation and they are not quantitatively evaluated. Instead, we treat the language-based description of errors made by computer vision models as the problem itself, providing means to evaluate them.
4 A family of approaches to solve LBEE
We propose a family of simple and efficient approaches to tackle the LBEE problem. In contrast with prior art [9, 51, 20], designed for classification and often requiring access to ground-truth or privileged information, our proposed methods are unsupervised and task agnostic. Our solutions rely on two key elements: i) a joint vision-and-language space $\bm{\mathcal{F}}$ , for which we use Open-CLIP [17], and ii) the sentence set ${\mathcal{S}}$ introduced in the previous section, assumed to be large enough to describe images from the target set and to contain potentially relevant reasons for model failure (see Appendix C).
We recall that the problem at hand is the following: Given a set of images, denoted by $\mathcal{X}$, a task-specific pretrained model $\mathcal{M}_{\theta}$, and a large sentence set ${\mathcal{S}}$, we want to select the subset from ${\mathcal{S}}$ that describes the samples from $\mathcal{X}$ on which the model $\mathcal{M}_{\theta}$ underperforms. Our proposed solution is illustrated in Fig. 1 and follows the steps detailed below.
Step 1: Splitting the target set into easy and hard subsets. Given the model ${\mathcal{M}}_{\theta}$ and a confidence measure $\varphi_{\theta}$, we define two thresholds $t_{\varphi}^{h}$ and $t_{\varphi}^{e}$ and split the target set ${\mathcal{X}}$ into three sets: an easy one ${\mathcal{X}}^{e}=\{{\bf x}_{i}\in{\mathcal{X}}|\varphi_{\theta}({\bf x}_{i})>t_{\varphi}^{e}\}$, a hard one ${\mathcal{X}}^{h}=\{{\bf x}_{i}\in{\mathcal{X}}|\varphi_{\theta}({\bf x}_{i})<t_{\varphi}^{h}\}$, and, if $t_{\varphi}^{e}>t_{\varphi}^{h}$, a neutral set with the other images. If ground-truth for the target set is available, we can use the model’s performance to split the data; if not, we can rely on any unsupervised confidence measure.
Step 2: Clustering easy and hard subsets. Let us denote by ${\mathcal{E}}_{\text{vis}}$ and ${\mathcal{E}}_{\text{txt}}$ the visual and text encoders mapping images and sentences into the joint space $\bm{\mathcal{F}}$ , respectively. Let $\Phi=\{{\mathcal{E}}_{\text{vis}}({\bf x}_{m})\}^{M}_{m=1}$ be the set of visual representations of images in ${\mathcal{X}}$ and let $\Phi^{h}$ and $\Phi^{e}$ be the set of representations associated with hard and easy samples, respectively. We cluster the samples in $\Phi^{h}$ and $\Phi^{e}$ independently, using an arbitrary clustering method [^1] We use Agglomerative Clustering with Ward-linkage [48] in our experiments. and represent each cluster by its centroid, referred also as its prototype. We compute it as the average of the embeddings of the samples in the cluster. Let ${\mathcal{C}}^{h}=\{{\bf c}^{h}_{i}\}_{i=1}^{|{\mathcal{C}}^{h}|}$ and ${\mathcal{C}}^{e}=\{{\bf c}^{e}_{j}\}_{j=1}^{|{\mathcal{C}}^{e}|}$ be the set of prototypes associated with hard and easy samples, respectively. In the remainder of the paper, we will use ${\bf c}^{h}_{i}$ and ${\bf c}^{e}_{j}$ to indistinctly refer to the clusters or to their prototypes.
Step 3. Matching sentences with prototypes. Let us denote by ${\bf s}_{n}={\mathcal{E}}_{\text{text}}(s_{n})$ the textual embeddings of the sentences $s_{n}$. These embeddings lie in the same joint embedding space as the images and, in turn, as the prototypes ${\bf c}^{e}_{i}$ and ${\bf c}^{h}_{i}$. Therefore, for each prototype ${\bf c}$ (easy or hard) we can compute the cosine similarity (i.e., the dot product between L2-normalized representations) to each textual embedding ${\bf s}_{n}$. We can store these similarities into a vector ${\bf v}\in\mathbb{R}^{N}$, where the $n^{\textrm{th}}$ element of this vector is $<{\bf c},{\bf s}_{n}>$. This allows us to characterize each cluster in terms of its semantic similarity with the sentences in ${\mathcal{S}}$.
Step 4: Retrieving sentences for hard prototypes. Here, we look for a set of peculiar sentences for each hard cluster. To select relevant sentences for a given cluster (easy or hard), we simply rank the sentences by their similarity to the cluster center (by ranking the elements of the vectors ${\bf v}^{e}_{j}$ and ${\bf v}^{h}_{i}$). Then we either retain the top ranked sentences or all sentences with a corresponding value in ${\bf v}$ above a predefined threshold $\tau$. We denote the set of retained sentences for ${\bf c}^{h}_{i}$ and ${\bf c}^{e}_{j}$ by ${\mathcal{S}}^{h}_{i}$ and ${\mathcal{S}}^{e}_{j}$, respectively. While ${\mathcal{S}}^{h}_{i}$ effectively describes images within ${\bf c}^{h}_{i}$, not all sentences will point to failure reasons. We need to find sentences describing the visual features that make the cluster hard. We propose therefore to contrast the hard cluster with its closest easy cluster: its proximity in $\bm{\mathcal{F}}$ implies content similarity, hence, contrasting allows isolating the attributes that characterize the hard cluster specifically, and not the easy one.
Let ${\bf c}_{j}^{e}\in{\mathcal{C}}^{e}$ be the closest easy prototype to ${\bf c}_{i}^{h}\in{\mathcal{C}}^{h}$ based on their distance in $\bm{\mathcal{F}}$ In the following, we propose three different methods to isolate sentences that are peculiar to the hard cluster ${\bf c}_{i}^{h}$. For a fair and more straightforward comparison between different methods, we fix the number of sentences retained for each cluster and for each method to a predefined $K$, i.e. $\lvert{\mathcal{R}}_{i}^{h}\rvert=K$, where ${\mathcal{R}}_{i}^{h}$ is the sentence set retained by a method for the hard cluster ${\bf c}_{i}^{h}$.
SetDiff: Sentence set differences. Given the pair of clusters (${\bf c}^{h}_{i}$,${\bf c}^{e}_{j}$), we first select relevant sentences for both—${\mathcal{S}}_{i}^{h}$ and ${\mathcal{S}}_{j}^{e}$—and then remove from ${\mathcal{S}}_{i}^{h}$ the sentences that are present in ${\mathcal{S}}^{e}_{j}$, yielding ${\mathcal{R}}^{h}_{i}={\mathcal{S}}^{h}_{i}\setminus{\mathcal{S}}_{j}^{e}$. Since this approach does not guarantee that $\lvert{\mathcal{R}}_{i}^{h}\rvert=K$, we fill ${\mathcal{S}}_{j}^{e}$ with sentences corresponding to values in ${\bf v}^{e}_{j}$ above a threshold [^2] We set $\tau=0.25$ based on the observation that the average similarity between our sentence sets and image sets is around 0.17 due to the well-known modality gap [25]. $\tau$ and build ${\mathcal{R}}^{h}_{i}$ with the $K$ most similar sentences to ${\bf c}^{h}_{i}$ that are not in ${\mathcal{S}}^{e}_{j}$.
PDiff: Prototype difference. Assuming that higher scores in the element-wise difference between the similarity vectors ${\bf d}^{h}_{i}={\bf v}^{h}_{i}-{\bf v}^{e}_{j}$ better describe the hard clusters than the easy ones [^3] This, up to some normalization, is equivalent to first consider the prototype differences and then compute the similarity between each sentence $s_{n}$ and ${\bf d}^{h}_{i}$., we can rank ${\bf d}^{h}_{i}$ and retain the sentences corresponding to the top $K$ values. We denote this set by ${\mathcal{S}}^{d}_{i}$ and define ${\mathcal{R}}^{h}_{i}={\mathcal{S}}^{d}_{i}$.
FPDiff: Filtered PDiff. A drawback of PDiff is that it does not guarantee that the selected sentences accurately describe the hard cluster. A remedy to this is to only retain sentences that have high similarity to the prototype ${\bf c}^{h}_{i}$, namely sentences in ${\mathcal{S}}^{d}_{i}$ that are also in ${\mathcal{S}}^{h}_{i}$. We define ${\mathcal{R}}^{h}_{i}={\mathcal{S}}^{h}_{i}\cap{\mathcal{S}}_{i}^{d}$. Also here, to guarantee that $\lvert{\mathcal{R}}_{i}^{h}\rvert=K$, we fill ${\mathcal{S}}_{i}^{h}$ with sentences corresponding to values in ${\bf v}^{h}_{i}$ above a threshold $\tau$ and build ${\mathcal{R}}^{h}_{i}$ with the top $K$ sentences having the highest values in ${\bf d}^{h}_{i}$ that are also in ${\mathcal{S}}^{h}_{i}$.
TopS. We also add a baseline called TopS , describing the hard clusters without any contrasting, by selecting the top $K$ ranked sentences from ${\mathcal{S}}^{h}_{i}$ (${\mathcal{R}}^{h}_{i}={\mathcal{S}}^{h}_{i}$).
Step 5: Producing the final output. The final set of sentences that describe potential reasons for failure for the whole dataset is the union of all the sentence sets retained for all the hard clusters, namely
${\mathcal{R}}_{{\mathcal{S}}}=\bigcup\nolimits_{i=1}^{|C^{h}|}{\mathcal{R}}^{h}_{i}\enspace.$ (2)
We conclude by positioning these methods with respect to prior works.
Relations with prior art. As mentioned in Sec. 3, previous work related to ours [9, 51] describes a problem formulation that does not perfectly align with ours. Yet, their approaches proposed to assign explanations to predicted error modes can be related to some of the methods we propose above. DOMINO [9], which associates sentences based on average CLIP embedding from which the average class representative is extracted, is closely related to PDiff : when we analyze the errors per class (see Sec. 6), this version of our method can be interpreted as a variant of DOMINO in which the class prototype is replaced with the closest easy prototype. FACTS [51] relies on a captioning tool [31] to assign a single tag to each partition without contrasting, hence, it is mostly related to TopS.
5 Evaluation metrics for LBEE
Our third contribution is a set of metrics to evaluate predicted language explanations by LBEE methods, which can retrieve error-related sentences without any supervision for an arbitrary visual task. Concretely, given the output of a method, namely the set of sentences ${\mathcal{R}}_{\mathcal{S}}\subset{\mathcal{S}}$, we need to assess i) whether these sentences actually point to reasons for model failure, ii) how well the retrieved sentences characterize the image set assigned to the cluster, and iii) how well the predicted sentence set ${\mathcal{R}}_{\mathcal{S}}$ covers the set of potential explanations given by ${\mathcal{S}}^{*}_{\beta}$. We propose metrics that allow measuring a method’s performance in these dimensions.
Our main goal is retrieving sentences $s_{n}$ that point to reasons for model failure, i.e. $s_{n}\in{\mathcal{S}}^{*}_{\beta}$. This means that the hardness score of the sentence $\omega_{\theta}^{s_{n}}$—formally defined in Eq. 8, Sec. C.2— satisfies the condition $\omega_{\theta}^{s_{n}}<\omega_{\theta}^{avg}-\beta$. The lower this value is, the more the sentence is correlated with model failure. We define the Hardness Ratio (HR) for a given cluster as the ratio of sentences among the retrieved ones in ${\mathcal{R}}_{i}^{h}$ (see Sec. 4) for which the above condition holds, and we define the AHR as the average across all clusters
$\text{HR}_{i}=\frac{\lvert{\mathcal{R}}_{i}^{h}\cap{\mathcal{S}}^{*}_{\beta}\rvert}{\lvert{\mathcal{R}}_{i}^{h}\rvert}\,\,\,\textrm{and}\,\,\,\text{AHR}=\frac{1}{|C^{h}|}\sum_{i=1}^{|C^{h}|}\text{HR}_{i}$ (3)
ii) Average Coverage Ratio (ACR). Next, we introduce a metric that assess the coverage of the sentences associated with the hard clusters, i.e. the ratio of images in the cluster for which the retrieved sentence holds. Intuitively this shows how well a sentence $s_{n}\in{\mathcal{R}}^{h}_{i}$ characterizes a hard cluster ${\bf c}^{h}_{i}$. $\text{CR}_{i}$ is the mean of these ratios taken over the sentences retained in ${\mathcal{R}}^{h}_{i}$. Formally,
$\text{CR}_{i}=\frac{1}{\lvert{\mathcal{R}}_{i}^{h}\rvert}\sum_{s_{n}\in{\mathcal{R}}_{i}^{h}}\frac{1}{\lvert{\bf c}^{h}_{i}\rvert}\sum_{{\bf x}_{m}\in{\bf c}^{h}_{i}}\Gamma({\bf x}_{m},s_{n})\enspace,$ (4)
where $\Gamma(.,.)$ is a binary operator that takes in input an image and a sentence, and outputs 1 if the sentence $s_{n}$ is relevant for the image ${\bf x}_{m}$ and zero otherwise (see Sec. C.2 for details). To obtain the final ACR value, we simply average $\text{CR}_{i}$ across all clusters, namely
$\text{ACR}=\frac{1}{|C^{h}|}\sum_{i=1}^{|C^{h}|}\text{CR}_{i}.$
iii) ${\mathcal{R}}_{\mathcal{S}}$ vs. ${\mathcal{S}}^{*}_{\beta}$. To assess if the union of the retained sentences ${\mathcal{R}}_{\mathcal{S}}$ (defined in Eq. 2) accurately covers the potential errors represented by ${\mathcal{S}}^{*}_{\beta}\subset{\mathcal{S}}$, we compute the True Positive Rate (TPR) of the retrieved sentences and the Jaccard Index (JI) between the two sets
$\text{TPR}=\frac{\lvert{\mathcal{R}}_{\mathcal{S}}\cap{\mathcal{S}}^{*}_{\beta}\rvert}{\lvert{\mathcal{S}}^{*}_{\beta}\rvert}\,\,\,\text{and}\,\,\,\text{JI}=\frac{\lvert{\mathcal{R}}_{\mathcal{S}}\cap{\mathcal{S}}^{*}_{\beta}\rvert}{\lvert{\mathcal{R}}_{\mathcal{S}}\cup{\mathcal{S}}^{*}_{\beta}\rvert}.$ (5)
TPR indicates how well a method covers the ground-truth explanations, while JI measures the overall coverage, also taking into account false positives.
Unsupervised NICO++{++}+ +          Per-class NICO++{++}+ +                  ImageNet                 Urban segmentation
Figure 2: Numerical results (all datasets). From top to bottom: ACR, AHR, TPR and JI scores on NICO++{++}+ +  unsupervised and supervised per-class (first and second column, respectively), ImageNet per-class (third column) and Urban Scene Segmentation (last column).
6 Experiments
Table 1: Examples of sentences from the user-defined sentence set ${\mathcal{S}}$. For each sentence we show if it belongs to ${\mathcal{S}}^{*}_{\beta}$ and to ${\mathcal{R}}_{\mathcal{S}}$ for TopS (TS), PDiff (PD), SetDiff (SD) and FPDiff (FP) for the datasets WD2, IDD and ACDC. “” means $s_{n}\in{\mathcal{S}}^{*}_{\beta}$ and “$\times$” means $s_{n}\notin{\mathcal{S}}^{*}_{\beta}$—namely, the hardness score of $s_{n}$ is below the required level. For the methods, “+”/“+” indicate that the sentence is in ${\mathcal{R}}_{\mathcal{S}}$, where “+” means true positive (i.e., the sentence is also in ${\mathcal{S}}^{*}_{\beta}$) while “+” means false positive. An empty space indicates that the sentence was not in the corresponding ${\mathcal{R}}_{\mathcal{S}}$.
WD2 IDD ACDC
Sentences (An image …) ${\mathcal{S}}^{*}_{\beta}$ TS PD SD FP ${\mathcal{S}}^{*}_{\beta}$ TS PD SD FP ${\mathcal{S}}^{*}_{\beta}$ TS PD SD FP
“taken at night” + + + + + + + +
“taken in the evening” + + + + + +
“taken at dusk” + + + + + +
“taken in a foggy weather” + $\times$ + + + +
“taken in a rainy weather” + + + $\times$ + + +
“taken in a snowy weather” + + + $\times$ $\times$ +
“taken in a dull weather” + + + + $\times$ $\times$ + + + +
“with reflection on the road” + + + + + +
“with shadows on the road” $\times$ + $\times$ + +
“with water on the road” + $\times$ $\times$ + +
“with mud on the road” $\times$ + + $\times$ + + $\times$
“with obstacle on the road” + + $\times$ + $\times$
“with road barrier on the road” $\times$ + $\times$ + + + + + +
“with rail track on the road” $\times$ + + + + + +
“with rocks on the road” $\times$ + + $\times$ $\times$
“showing a highway scene” $\times$ + $\times$ + + $\times$ + + + +
“showing an industrial scene” $\times$
“showing construction site” $\times$ + + + $\times$ + +
“showing sub-urban scene” $\times$ + + $\times$ + +
“with rickshaw on the road” $\times$ + + + + $\times$ + + + +
“with motorbike on the road” $\times$ + + $\times$ + +
“with vehicle on the road” $\times$ + $\times$ + $\times$ + + + +
“with tram on the road” $\times$ $\times$ + $\times$ + + +
“with jeep on the road” $\times$ + + + $\times$ $\times$
“with animal on the road” $\times$ + + + + $\times$
“with people on the road” $\times$ + + $\times$ + + $\times$ +
“with crowded foreground” $\times$ $\times$ + + + $\times$
“with motion blur” + + + + + +
“with underexposure” $\times$ + + $\times$ +
“with low contrast” $\times$ + + $\times$ +
“of a tunnel” + + + + + + + + +
“of fences” $\times$ $\times$ + +
“of guard-rail” $\times$ $\times$ $\times$ + + +
“of a traffic jam” + + + $\times$
“of traffic lights” $\times$ +
“of a parking” $\times$ + + + + $\times$ +
Experimental setup. We tackle three tasks: semantic segmentation of urban scenes, classification under the presence of spuriously correlated data, and large-scale ImageNet classification. For the first, we consider a ConvNeXt [29] segmentation model trained on Cityscapes [3] and test on three challenging datasets: WD2 [52], IDD [45] and ACDC [37]. For the second task we train ResNet50 image classification models on different spuriously correlated data derived from NICO++{++}+ + and demonstrate how the proposed approaches can be used for understanding model failures due to such subtle biases [55, 14]. We consider two cases, an unsupervised case where the full set is split with class prediction entropy and a supervised one, where following [9, 51] we analyze the performance on each class separately. In the latter case we use the class prediction to split the data into easy and hard sets and, in addition to evaluating LBEE, we compare our simple split-and-cluster partitioning to the more complex partitioning methods from [9, 51], using their evaluation protocol (see in Appendix B). Finally, to showcase the scalability of our approaches, we analyze different pre-trained architectures on the ImageNet 1K validation set, focusing on the top-1 classification performance and seeking for explanations for each class individually. We provide details about each task, related models and the datasets in Appendix A.
Default design choices. We use Open-CLIP [17] as vision-and-language embedding space. To generate the GT sentence set ${\mathcal{S}}^{*}_{\beta}$, we use $\beta=.2*\omega_{\theta}^{\text{std}}$ with $\omega_{\theta}^{\text{std}}$ being the standard deviation of $\omega_{\theta}$ computed over ${\mathcal{X}}$. We set the number of hard clusters to $C=15$ for large datasets and $C=5$ for the per-class data analyses. We use the same number of clusters for the easy set ($|calC^{h}|=|{\mathcal{C}}^{e}|=C$). To split the data we use the output entropy in the unsupervised cases and class probability or ranking in the per-class case (see details in Appendix D). We retain three sentences for each cluster($K=3$).
In the following, we show numerical results and illustrate them with some qualitative examples. We report in-depth analyses for the different datasets, also concerning the sensitivity to crucial hyper-parameters.
TopS SetDiff PDiff FPDiff
"taken in a snowy weather"    "with underexposure"    "taken in a foggy weather" "taken in a foggy weather"
"taken in a cloudy weather"    "showing a city scene"    "taken in a windy weather"    "taken in a windy weather"
"taken in a dull weather"    "with low contrast"   "taken in a dull weather"    "taken in a dull weather"
TopS    SetDiff    PDiff    FPDiff
taken in a rainy weather    “shadows on the road    “water on the road    “water on the road
taken in a dull weather    “people on the road    “taken in a rainy weather    “taken in a rainy weather
taken in a cloudy weather    “of a sidewalk    “rickshaw on the road    “taken in a stormy weather
Figure 3: Two hard clusters in ACDC explained by different methods. We illustrates a few example images from the cluster and below the three sentences retained by different methods for this cluster.
Figure 4: Unsupervised and supervised splitting (ACDC). Results of different metrics (from left to right ACR, AHR, TPR and JI) on the ACDC dataset when the split is done with entropy, pixel accuracy and the mIoU.
6.1 Results
In Figs. 3, 5, 7, 9 and 11 we report qualitative results, showing the sentences selected by our methods for various hard clusters. In all cases, each row represents a hard cluster, and below we report the three sentences provided by the different methods.
Focusing, e.g., on Figs. 3, 5 and 7 (urban segmentation on ACDC, IDD and WD2, respectively) we can see that our methods can successfully isolate the main reasons for clusters to be hard, for example difficult visual conditions (rainy/foggy weather or image taken at night) or the presence of elements unseen during training such as tunnel,mud or light reflections). Indeed, such elements can represent real challenges for a model trained on Cityscapes as observed also e.g. in [4].
With our methods, the user can easily asses the potential failures at a glance, instead of inspecting the hundreds of images contained in the above clusters [^4] The user could rank easily the images in the cluster for each selected sentence for the visual inspection based on the corresponding Open-CLIP similarity scores.. Such information can be further used for different down-stream tasks, e.g., helping the user to select which new samples to annotate [41, 33].
In Fig. 2, we provide numerical results for the different methods, tasks and datasets, using the metrics introduced in Sec. 5 and the default setting and hyper-parameters detailed above. First, we observe that TopS  is the best performing method in terms of ACR scores. This is not surprising, since sentences closest to the prototype have obviously the highest coverage. Yet, these sentences are not necessarily pointing to error explanations—indeed, see the low AHR scores in some cases, with IDD being the most dramatic one (right-most column, third panel).
Concerning the metrics related to the failure explanations (AHR/TPR/JI), in both NICO++{++}+ +  scenarios, where we have ground truth relevance scores, and on WD2 FPDiff performs best or close to it. It performs worse on ACDC and IDD: there is no clear winner among the proposed methods in these two datasets, since the higher performance of PDiff  and SetDiff  in terms of TPR and JI comes at the price of a lower ACR.
In the following, we provide more in-depth analysis focusing on the different datasets.
TopS SetDiff PDiff FPDiff
"rickshaw on the road"      "shadows on the road"      "dirt on the road"      "dirt on the road"
"construction on the road"      "residential scene"      "countryside scene"      "construction on the road"
"sub-urban scene"      "debris on the road"     "mud on the road"     "debris on the road "
TopS SetDiff PDiff FPDiff
traffic    “shadows on the road    “highway scene    “highway scene
rickshaw on the road    “crowded background    “traffic jam    “traffic jam
showing an sub-urban scene    “light reflections on the road    “advertising board    “traffic
Figure 5: Two hard clusters from IDD explained by different methods. We illustrates a few example images from the cluster and below the three sentences retained by different methods for this cluster.
Figure 6: From left to right, we show how the ACR, AHR, TPR and JI metrics vary on the IDD experiment, by varying the number of sentences selected for each cluster $K$ (top) and varying the number of clusters $C$ (bottom). Note that $C=1$ corresponds to not performing any clustering and, hence, analyzing the whole hard set versus  the whole easy set. The very poor numbers obtained with $C=1$ emphasizes the importance of clustering.
6.1.1 ACDC
To recap, we test models trained on Cityscapes (daylight and clear weather) on ACDC, which contains samples recorded under various weather and daylight conditions. Results in Fig. 2 show that the best AHR performance is obtained with TopS , but also that this method is the worst performing one in terms of TPR and JI. For what concerns the latter metrics, PDiff  shines as the best performing method. To shed light on these results, we complement them with Tab. 1 (and with the more detailed Tab. 5) showing sentences from the User-defined sentence set ${\mathcal{S}}$ that are either in ${\mathcal{S}}^{*}_{\beta}$ or in any ${\mathcal{R}}_{\mathcal{S}}$, namely the ground truth set of hard sentences and the ones retrieved by our methods, respectively.
From these results we can make the following observations. i) The main advantage of TopS  comes from retrieving the sentences related to images taken at night or in the evening. ii) The other methods fail in retrieving these sentences because the closest easy clusters also contain night or evening images. iii) On the other hand, these methods often select sentences referring to “fog”, or “rain” (see e.g. Fig. 3) that are indeed valid characteristics of the images in the cluster, but yet these conditions do not affect sufficiently the model to have a hardness score that satisfies the condition $\omega_{\theta}^{s_{n}}<\omega_{\theta}^{\text{avg}}-\beta$.
Varying the easy/hard splitting strategy. In order to evaluate the effectiveness of the entropy as a metric to split datasets into easy and hard partitions, in Fig. 4 we compare the results obtained by splitting the ACDC set with entropy versus  using metrics that rely on annotated samples, namely the pixel accuracy (average of correctly predicted pixels) or mIoU.
We can observe that while using the GT mIoU improved significantly the ACR and AHR scores (averaged over all clusters), globally the TPR and JI changed only slightly. This suggests that, in general, the sentences retained by PDiff , FPDiff and SetDiff  have better coverage in the cluster (higher ACR) and retain more often a hardness score that satisfies the required constraint. The TPR and JI scores are comparable across splitting strategy. By analyzing the sentences retrieved in each setting, we discovered that with GT mIoU the contrastive methods are able to recover sentences such as “taken at night”, “taken in the evening” and “taken at dusk”, but fail in retrieving sentences such as “with road barrier”/“with rail track on the road” and “with motion blur“/”with underexposure”, sentences that were retrieved by using the entropy.
Finally, sentences referring to fog and rain, having scores slightly below the required condition $\omega_{\theta}^{s_{n}}<\omega_{\theta}^{\text{avg}}-\beta$, are present in ${\mathcal{R}}_{\mathcal{S}}$ for all splitting strategies and most methods. One could interpret those as less severe false positives.
6.1.2 IDD
In the case of IDD, the best AHR/TPR/JI results in Fig. 2 are obtained with SetDiff . Yet, the lower ACR suggests that the sentences retained by this method have lower coverage in the cluster than the ones retained by the other methods. IDD is a large dataset with high content variation, hence, posing more challenges. Indeed, splitting into 15 clusters makes the content in each cluster very heterogeneous (see Fig. 5) and therefore the selection of the three most representative sentences is rather difficult.
To tackle this complexity, we carry out two additional IDD experiments, varying i) the number of sentences selected and ii) the number of clusters.
6.1.3 WD2
GT-based UserDef LLM-based
wheeled vehicle      “mud on the road      “off-road vehicle
trailer     “rocks on the road     “taken on a mountainous road
road     “jeep on the road     “taken on a rocky road
GT-based UserDef LLM-based
streetlight      “taken at night      “taken in a vibrant night
traffic light     “taken at dusk     “taken in the late snowy night
dashcammount     “jeep on the road     “taken in a gloomy night
GT-based UserDef LLM-based
tunnel      “tunnel      “tunnel
streetlight     “light reflection on the road     “underwater tunnel
dashcammount     “motion blur     “underpass
Figure 7: Three out of 15 hard clusters from WD2 (unsupervised case) explained by FPDiff  when we use different the sentence sets (GT based, User defined or LLM generated).
Figure 8: Varying the sentence set. Results obtained on WD2 when ${\mathcal{S}}$ is the GT-based, User-defined and the LLM-generated sentence set (see Sec. C.1).
Varying the number of retained sentences ($K$). We show in Fig. 6 (top) results for IDD as we vary the number of retained sentences per cluster. We observe that with the increase of the number of sentences retained, the ACR increases for FPDiff , but decreases for the other methods; on the other hand, the ACR is stable starting from $K=5$. The AHR is generally stable as we vary $K$, except for FPDiff , where the best value is obtained with $K=1$. Both TPR and JI benefit from the increase of the number of sentences (higher $K$). Overall, we can confirm that SetDiff  outperforms the other methods in terms of AHR, TPR and JI. This result, combined with the low ACR, suggests that SetDiff  is more capable of detecting rarer failure reasons, for which there are less support example images in the cluster/dataset.
Varying the number of clusters ($C$). In Fig. 6 (bottom) we present results obtained on IDD when we vary the number of clusters ($|{\mathcal{C}}^{h}|=|{\mathcal{C}}^{e}|=C$). By increasing the number of clusters, we observe an increase of the TPR (as expected) for all methods—with PDiff  being the one that benefits the most. Concerning the other metrics, the default value $C=15$ is a reasonable trade off. These experiments suggest that IDD is overall a very challenging set, with some of the failure causes under-represented—especially the ones related to the weather conditions (see sentences in Tab. 1 and Tab. 5), or not clustered together according to these characteristics—cf. Fig. 5. Furthermore, note that as the filtering in FPDiff relies on the TopS  ranking, hence, the very poor performance of TopS  on this particular set (IDD) affects negatively the performance of FPDiff , making PDiff  perform better than FPDiff when we consider more clusters or retain more sentences. Finally, the poor performance obtained with $C=1$—corresponding to running the methods without the clustering step—show that describing the full hard set by contrasting with the full easy set is a sub-optimal solution.
TopS SetDiff PDiff FPDiff
taken in autumn       “taken in autumn       “sailboat in the water    “sailboat in the water
with waterways   “with waterways       “sailboat in the rocks       “sailboat
plants in the water   “plants in the water   “goose in the water   “birds in the water
TopS   SetDiff   PDiff   FPDiff
train       “train       “ train in the rocks       “ train in the rocks
train in the rocks   “train in the rocks       “bus in the rocks       “bus in the rocks
bus in the rocks   “bus in the rocks   “train in the grass   “train in the grass
TopS   SetDiff   PDiff   FPDiff
truck taken at sunset       “truck taken at sunset       “truck taken at sunset       “truck taken at sunset
bus taken at sunset   “bus taken at sunset       “bus taken at sunset       “bus taken at sunset
train taken at sunset   “train taken at sunset   “truck in the water   “truck in the water
Figure 9: Three out of 15 hard clusters from NICO++{++}+ + 85 (unsupervised case) explained by different methods.
Sensitivity to the choice of the sentence set. We analyse here the performance of the different methods as we vary the sentence set ${\mathcal{S}}$. For this experiment, we use the WD2 dataset, because for this dataset we have both User-defined and GT-based sentence sets. We also compare with the significantly larger LLM-based sentence set (see details in Appendix C). We show qualitative comparison in Fig. 7 and numerical results in Fig. 8.
First, we notice that both the User-defined and the LLM-generated sentence sets yield similar ACR and AHR scores. Yet, for the latter we have many sentences that are semantically related and carry similar hardness scores, for example “taken at night”/“mesmerizing moonlight”/“spellbound starlight”/“enigmatic night”/“miraculous midnight”. Since we limit the number of sentences each method can retrieve, in this case we cannot retrieve all of them with a small $K$ and, in turn, this will result into significantly lower TPR and JI scores.
When we use the GT-based sentence set with $o=0.2$, ${\mathcal{S}}^{*}_{\beta}$ contains only five sentences that are “showing a tunnel”, “showing a bridge”, “showing traffic light”, “showing a bus” and “showing snow”. Here, all methods introduce many false positives (explaining the low JI score). PDiff  and SetDiff  were able to retain four out of five sentences (except “showing a bus”) and TopS  and FPDiff  three out of five (missing also “showing a bridge”). The missing sentences had not enough image coverage in the clusters. The AHR and ACR scores are also low, due to the fact that it is not easy to find shared content for all clusters (see the low ACR even for TopS ). Finally, note that the GT-based sentence set mainly describes the presence of the classes (“Image showing <class>"), therefore, it is not sufficient to describe the hard clusters to the desired level of details. This highlights the importance of properly designing ${\mathcal{S}}$, as well as the importance of using additional text resources to analyze the performance of our models.
Figure 10: Varying $\beta$ (by varying $o$). Results obtained on NICO++{++}+ + 85 (unsupervised). The value $\beta$ defines the ground truth set ${\mathcal{S}}^{*}$, so its value impacts only the AHR/TPR/JI metrics (not the output of the methods or ACR). On x-axis: we plot $o$, where $\beta=o\cdot\omega_{\theta}^{\text{std}}$ with $\omega_{\theta}^{\text{std}}$ being the standard deviation of $\omega_{\theta}$ (entropy) computed over ${\mathcal{X}}$.
TopS SetDiff PDiff FPDiff
stage indoor       “basement       “with motion blur       “blurry image
arena performance       “stage indoor       “blurry image       “stage outdoor
taken in a basement       “thumbnail image   “stage outdoor   “stage indoor
TopS SetDiff PDiff FPDiff
acoustic guitar       “taken in recreation room       “in dorm room       “taken in motel
electric guitar       “taken in motel       “market indoor       “recreation room
banjo       “taken in shed   “taken in storage room   “music studio
TopS SetDiff PDiff FPDiff
taken in dressing room       “with reflections       “ taken in promenade       “taken in promenade
taken in house       “with overexposure       “ taken in street       “ taken in street
blurry image       “screen shot   “stage outdoor   “stage outdoor
Figure 11: Analyzing EfficientNet performance on the ImageNet 1K. The first two rows are from the “acoustic guitar” class, and the third is from the bath tube class.
6.1.4 NICO++{++}+ + 85
We provide in LABEL:{fig:qualitativeNPPmain} qualitative examples for NICO++{++}+ + 85, in the unsupervised case (where we use the prediction entropy to split the data). We observe that our methods can capture the “context-class” associations that were rarely seen in the training set [^5] See in Tab. 2 the distribution of the images in the validation set over the different classes-context pairs. Note that the validation set follows the training distribution. and hence the model struggles with their classification. Indeed we find “birds” or “plants” in “water” (top row), “landways” in the “rocks” (middle row) and “landways” in ‘dim lighting” (bottom row).
Next, we provide a sensitivity study on this dataset when we vary the hardness threshold value $\beta$.
Varying $\beta$ (by varying $o$). Recall that $\beta$ is the hyper-parameter used to select the ground-truth sentence set ${\mathcal{S}}^{*}$, where higher $\beta$ values imply retaining only the sentences with the highest hardness scores as valid failure descriptions, while lower values imply that sentences with a score $\omega_{\theta}^{s_{n}}$ below $\omega_{\theta}^{\text{avg}}$ but with a smaller margin $\beta$ are also considered as valid explanations. In Fig. 10 we show the impact of varying $\beta$ on the AHR, TPR, and JI metrics. We carried out this experiment on the spurious classification with NICO++{++}+ + 85, considering the unsupervised case where we have GT image-sentence relevance scores (top row). As the number of sentences retained by the method is fixed ($K=3$), varying $\beta$ only affects ${\mathcal{S}}^{*}$ and not ${\mathcal{R}}_{\mathcal{S}}$. Therefore, the value of ACR (that does not depend on $\beta$) is the same for each $o$ value. Intuitively, with this study we assess how much each method focuses on the hardest sentences. Indeed, increasing the value of $o$ (and hence increasing $\beta$) corresponds to accepting fewer and fewer sentences as valid explanations, namely the ones with the highest hardness score (lowest average accuracy). We observe that the FPDiff method, while also experiencing a drop in performance as we make the problem harder, holds a better performance than the others methods for all the values of $o$.
6.1.5 ImageNet 1K
Finally, we focus on ImageNet 1K, where we treat each class separately. We can make several observations. i) When the error slices (hard clusters) mainly contain false positives from semantically similar classes or false negatives with shared content, TopS will generally capture the shared characteristics—see examples in Fig. 11 (top two rows), Fig. 13 and Fig. 14 in Appendix E. ii) In the case of false positives with semantically similar classes, e.g.electric guitar” and “banjo” in place of the “acoustic guitar” class (see first row in Fig. 11), the main difficulty for the other methods is that the same sentences are also highly ranked for the easy clusters, making it difficult for them to correctly identify the desired explanations—namely, contrasting the explanation between the easy and the hard cluster. This limitation is inherited from CLIP, which has difficulties to distinguish fine-grain classes. iii) When the false negatives share some semantic similarities and there is a candidate sentence that well describes this, all methods are able to find it. For example, see in Fig. 11 (second row) the difficulty of the model to label the “acoustic guitar” when the photo is showing the person holding the guitar on a “stage”. Note that for these images PDiff and FPDiff also identified the blur as potential challenge (or “low contrast” and “dim weather” condition, in Fig. 13). iv) Finally, we have clusters where even for a human is difficult to describe what is common in the images beyond the object class. When there is a mix between false positives and false negatives, all methods including TopS have difficulties to identify the failure reason. In these cases, the selected sentences are relevant to some of the images in the cluster but they do not point to the reason for the failure (see Fig. 11 last row).
7 Concluding remarks
Assessing model performance, and in particular failure modes, for arbitrary task in arbitrary environments should be easy, efficient and interpretable. While the gold standard for performance evaluation is using a human-annotated test set, this process is costly and cannot scale for computer vision models to be deployed in many and diverse scenarios. Moreover, quantitative evaluation measures such as dataset-level accuracy values do not fully reflect the details of the model performance. It is important to know what lies beyond performance values and if the model is failing on particularly critical images.
In this work, we posit that it is of the utmost importance being able to assess model performance on non-annotated samples, and to move beyond hard-to-interpret numbers. We address this by first formulating the LBEE problem and designing a family of task-agnostic approaches that do not require error specifications or user-provided annotations. We complement these two contributions with different metrics to benchmark methods for LBEE and rely on those to carry out in-depth analyses. We show that we can retrieve relevant sentences pointing to important errors, e.g., related to environments beyond our model’s comfort zone.
In the future, we plan to overcome the need to provide the sentence set ${\mathcal{S}}$, a design choice that makes evaluation more tractable, but that also limits our methods. Defining the sentence set ${\mathcal{S}}$ a priori has the advantage that our methods can be deployed off the shelf without any overhead, but yet the set might omit certain unexpected failure reasons. A possible solution could be to build the sentence set ${\mathcal{S}}$ directly from the target set ${\mathcal{X}}$, e.g., by running a captioner to describe all its images and merge these descriptions into the sentence set. Yet, in addition to being prohibitively expensive, a solution like this needs to be iterated for every new target set before running any method. Furthermore, in the case of a large image set, it would also require further steps to remove redundancy and to regroup similar sentences.
Appendix
Appendix A Datasets and Tasks
In the following, we detail the three main tasks we tackle in this paper: semantic segmentation of urban scenes, classification in the presence of spuriously correlated data, and ImageNet-1K classification.
Urban scene segmentation. We consider a ConvNeXt [29] segmentation model trained on Cityscapes [3], comprised of images from 50 European cities collected at daytime in clear weather conditions. We build our evaluation sets ${\mathcal{X}}$ using three datasets: i) WildDash 2 (WD2) [52], which contains challenging visual conditions such as motion blur, various road types, difficult weather, etc.  and ii) India Driving Dataset (IDD) [45], which contains images from Hyderabad, Bangalore and their peripheries, iii) ACDC [37] contains images recorded in adverse conditions, namely fog, rain, snow and night representing challenging domain shift for the Cityscapes model. For the main experiments we use the user-defined sentence set $\mathcal{S}$ with 148 sentences describing content related to urban scenes, road conditions, and image quality (described in Appendix C). In Appendix D, we show further results on WD2 with both a smaller GT-based sentence set derived from the metadata information available with the dataset and a larger sentence set of 1016 sentences generated automatically by an LLM (both detailed in Appendix C).
[…]
To evaluate error descriptions of the hard clusters, we build a sentence set $\mathcal{S}$ with 130 sentences that include information related to the the six classes, the six contexts, various sub-classes and their combination. Note that in this case we have GT relevance $\Gamma({\bf x}_{m},s_{n})$ between sentences and images.
We consider two cases, an unsupervised and a supervised one. In the former case we consider the full set and split it with entropy values computed on class predictions. In the latter case, similarly to [9, 51], we consider each class separately and use the class probability scores to split the data into easy and hard sets.
ImageNet-1K classification. We consider three architectures (ResNet-50 [12], VIT-B-16 [7], and EfficientNet-B1 [43]) trained on ImageNet-1K [36] and evaluate these models on the validation set. We build ${\mathcal{S}}$ as the union of a set of sentences {“An image of a <class>”} (for the 1K classes) with a set of sentences corresponding to place classes from Places365 [56] {“An image taken at <location>”}, and a set of sentences describing image type (photo, drawing), image quality (blurry, noise, JPEG compression), weather conditions, etc.  In total we have 1417 sentences (see details in Appendix C). To assess the relevance between a sentence and an image, we use a combination of ground truth based information (the ones related to the 1000 class names) and a VQA model to assess the relevance for the other sentences (see Sec. C.3).
In particular, we focus on the top-1 classification performance of ImageNet models and aim at analyzing failure cases for this task—namely, the reasons for not predicting the correct class on top. In ImageNet classification, several errors arise from confusions between similar classes. For example, the model will miss-classify dog breeds, bird/fish species, similar objects, etc. In such cases, knowing the correct class label can help to deduce the main reason from the retrieved description. Indeed, for example, if the selected sentence is {“An image of a water snake”} without the knowledge of the correct class, it is difficult to interpret whether the model predicts false positives (images of different type of snakes) or the model failed to recognize the water snake (false negative) for some reasons. Furthermore, while, class independent failure reasons such as those related to image capturing conditions (e.g.blurred or low contrast images) might emerge when processing the full dataset as a whole, errors related to the image content is more difficult to interpret if it is a failure reason without the target class (e.g.finding that the image was taken at a beach has different effect on the model for the swimming suit class than for the cow or ski suit class). Finally, even if we use a very large number of clusters with the full set, the content of the nearest easy cluster might differ in many aspects from the hard one, making it difficult to select the actual failure reason.
For all these reasons we opted to follow the prior art and apply LBEE methods by analyzing the data for each class independently. Still, in contrast to prior art, where only false negative errors are considered (the images that belong to the given class but labeled with a wrong class label), we also include false positives. The easy set ${\mathcal{X}}^{e}$ contains the images where the class has been correctly predicted (images from the current class).
Appendix B Comparison with prior art
DOMINO [9] and FACTS [51] mainly focus on evaluating the error mode discovery [9, 51], assuming to know the set of error types, and limit the language description of these error modes to a qualitative role. In contrast, in our problem formulation, we treat the language-based description of errors made by computer vision models as the problem itself. For this reason, a direct comparison with prior art is not straightforward. In this section, we make an attempt by simply comparing our split-and-cluster partitioning to the more complex, learnt partitioning methods by [9, 51], following their evaluation protocol.
[…]
To tackle this setting, [51] propose to learn a partitioning Gaussian Mixture model on a Biased validation set that follows the same bias as the training set and evaluate the partitioning on an Unbiased set where all class-context are equally represented (see statistics in Tab. 2). Note that this is done for each class independently (per-class case). Using the code provided by [51] [^7] https://github.com/yvsriram/FACTS, we learn and evaluate all possible set combinations (see Tab. 3). Note that (Biased-Biased) and (Unbiased-Unbiased) sets means that we learnt and test the partitioning function on the exact same data.
In our case, there is no learning involved: we simply split and cluster the data as discussed in Sec. 4. Then, for a given ($A,B$) pair of sets, where $A,B\in\{{\textit{Biased}},{\textit{Unbiased}}\}$, we assign images from the set $B$ to the cluster prototypes obtained with the images in set $A$, where the union of easy and hard prototypes is considered. We use the class probability to split the data—the same measure used by FACT to learn the partitioning function.
Evaluation protocol. Before discussing the comparative results, we recall the metrics and the evaluation protocol we use, proposed by [9, 51]. In particular they propose to rely on the Precision-at-K (P and K ) to measure how accurately a partitioning method aggregates samples from the same context given a class. Formally, let ${\mathcal{X}}$ denote the test set, $\{{\bf z}_{n}\}_{n=1}^{N}$ the ground-truth partitions and $\{\widehat{{\bf z}}_{m}\}_{m=1}^{M}$ the predicted set of discovered partitions. Given $K$, for each ground-truth partition ${\bf z}_{n}$ the most "similar" predicted partition $\widehat{{\bf z}}_{m}$ is selected based on the size of the intersection ${\bf z}_{n}\cap\widehat{{\bf z}}^{K}_{m}$, where $\widehat{{\bf z}}^{K}_{m}$ is the subset containing the top-$K$ elements of ${\mathcal{X}}$ according to the likelihood of belonging to $\widehat{{\bf z}}_{m}$. For FACTS/DOMINO the ranking corresponds to the partitioning function’s output, while in our case to the distance to the prototype. If we denote by $\widehat{{\bf z}}^{K}_{nm}$ the set that has the largest overlap with ${\bf z}_{n}$ given $K$, the P and K is defined as
[…]
where the average is computed, as in [51], over the five bias-conflicting groups (excluding the spuriously correlated GT partitions, such as waterways in water).
Table 2: The number of images in the Biased validation set for each (class, context) pair in corresponding three NICO++{++}+ + datasets. The Unbiased validation sets have 50 images for each (class, context) pairs in each of the three sets. The spurious contexts (1 to 6) are in order rock, grass, dim lighting, autumn, water, outdoor.
NICO++{++}+ + set
class $\downarrow$ context $\rightarrow$
mammals
birds
plants
airways
waterways
landways
[…]
1 2 3 4 5 6
638 4 2 2 3 3
7 321 2 2 3 3
7 4 154 2 3 3
7 4 2 220 3 3
7 4 2 2 266 3
7 4 2 2 3 277
[…]
1 2 3 4 5 6
638 4 2 2 3 3
7 321 2 2 3 3
7 4 154 2 3 3
7 4 2 220 3 3
7 4 2 2 266 3
7 4 2 2 3 277
[…]
1 2 3 4 5 6
638 10 5 7 8 9
19 321 5 7 8 9
19 10 154 7 8 9
19 10 5 220 8 9
19 10 5 7 266 9
19 10 5 7 8 277
Experimental results. We show in Tab. 3 the P and K values averaged over the six classes for the different NICO++{++}+ + sets. Each time we show results averaged over three different variants of the current model [^8] Since we were not able to reproduce the original paper’s numbers with the code, namely, the results corresponding to the (Biased,Unbiased) pairs, for fairness we ran their code with three different seeds and provide here the averaged results. We do not show the variances in the table to keep it simple, but the values are most often between 1 and 2.. In our case, we used the exact same ResNet50 models and the same Biased and Unbiased sets on which we perform the split and the clustering. We use five easy and five hard clusters as it is our default setting for the per-class case and as in the code from [51] the default number of slices was also set to ten.
From Tab. 3 we can see that our split-and-cluster method provides competitive, or sometimes even better, slice discovery performance than the more complex partition learning methods. Our method performs slightly worse than FACTS on the Unbiased set both when their partitioning was trained on the Unbiased or on the Biased set. Surprisingly though, both FACTS and DOMINO underperform when tested on the Biased set even when the partitioning function was trained on this set. In contrast, our method is more robust and performs similarly on both sets. When the FACTS model is trained on the Biased set, which strongly relies on priors about class-context associations (the configuration shown in their paper), it outperforms our model, for which the P and K decreases as the amount of class-context correlation increases (bottom part of Tab. 3). We hypothesize that the reason behind this result is that we do not have enough data for a proper clustering (see Tab. 2), while FACTS manages to learn the partitioning function even with a a few samples by exploiting the strong correlation between the classes and dominant context. As the model is trained mainly with images with specific combinations of class and correlated context—for example, mammals in rocky environments—the model rarely sees images from other classes in this context. Consequently, when the model learns to predict a class $c$" with high probability, it also associates this high probability with the presence of the corresponding context (e.g.rocks) even if the class (mammals) is absent. Therefore, when analyzing another class, such as plants, and learn the partitioning based on all six probabilities, the partitioning function primarily learns to assigns images based on the highest probability score, which is an indicator of the context. For instance, a plant image taken in a rocky environment will have a high probability score for the mammal class, indicating the presence of rocks. Note that in our case, when analyzing the plant class, we disregard the probabilities of other classes, thereby using much less prior information about the image context.
In summary, despite the simplicity of our method, it performs favorably to prior art in their proposed settings, while also being able to provide natural language descriptions of error modes as shown in Sec. 6.
Table 3: Comparisons with prior art. We compare the clustering obtained via Step 3 of our proposed methods (see Fig. 1) with the Domino and FACTS using averaged P and K scores with $K=10$. For our method, we split the dataset into 5 “easy” and 5 “hard” clusters and use their union as final partitioning. On the top rows (below Random) we show results obtained with the partitioning learnt on the Unbiased set and on the bottom rows the results were obtained when the partitioning is learnt on the Biased set. Random corresponds to assigning the test data randomly to the partitions.
[…]
         Test set $\rightarrow$ Unbiased Biased Unbiased Biased Unbiased Biased
Random $29.7$ $12.6$ $32.0$ $14.0$ $33.0$ $14.6$
Domino [9] Unbiased $44.4$ $12.1$ $47.0$ $12.3$ $52.7$ $22.3$
FACTS [51] 58.8 $28.1$ 63.6 $30.0$ 68.9 $39.9$
Ours(5-5) $53.6$ 52.3 $54.2$ 58.4 $55.3$ 66.1
Domino [9] Biased $26.0$ $17.1$ $19.0$ $19.8$ $17.4$ $25.6$
FACTS [51] 53.3 $31.4$ 53.2 $28.7$ 58.9 $34.4$
Ours(5-5) $51.3$ 52.2 $46.8$ 52.5 $41.9$ 46.5
Appendix C Defining the sentence sets
Our paper focuses on two objectives, 1) identifying error modes in computer vision models and 2) describing them with natural language. To achieve this, we rely on a large set of sentences (${\mathcal{S}}$), that can either be provided or automatically generated. This sentence set is expected to cover various potential reasons for model failures given a task. In this section, we first discuss three approaches to construct such sentence sets (Sec. C.1); then we detail how we create the ground-truth sentence set ${\mathcal{S}}^{*}_{\beta}$ for LBEE (Sec. C.2); finally, we show how to obtain pseudo-GT relevance scores between images and sentences using Visual-Question Answering (VQA) tools (Sec. C.3).
C.1 Building the sentence set ${\mathcal{S}}$
There are several possibilities to construct the sentence sets ${\mathcal{S}}$. We describe a few possibilities below.
User-defined. The set of sentences can be designed ad hoc for the task at hand by the end user who has the expert knowledge and can envisage the main difficulties a model might face. While such sentence set might be non-exhaustive, it can cover the most important failure cases. In this case, our methods are able to select the ones that confirm the user’s concerns (with additional visual support).
We follow this setting to build ${\mathcal{S}}$ in our experiments on the urban scene segmentation task. More specifically, we manually defined 148 sentences describing content related to urban scenes (buildings, pedestrians, traffic, two-wheels, trees, animals, garbage, etc. ) in the form of {“An image of a <urban scene content>”}, road conditions (weather, lighting, season) in form of {“An image taken in/at <condition>”} or image quality (underexposure, motion blur, etc. ) in form of {“An image taken in/at <effect>”}. A large subset of this sentence set can be seen in Tab. 5.
In this case, to get pseudo-ground-truth relevance between images and a sentence, we resort to VQA (as described in Sec. C.3).
GT-based. If available, we can use ground-truth information (class labels and metadata) to generate a sentence set. This is quite limiting in general: we mainly use this option because it yields the advantage of providing the GT relevance between images and sentences. We used this option in all NICO++{++}+ + experiments, where we build 130 sentences that include combinations from the six classes, the six contexts, and various sub-class information. For example, we have {“A photo of <sub-class> in the <context>”} such as “A photo of a cactus in the water”.
In the case of ImageNet-1K [36], the first 1K sentences (out of 1417) were also generated from the GT information, namely the class names. We consider sentences with the form {"An image of a <class>”} to which we added sentences with the form {“An image taken in <location>”} filling <location> with the place class names from Places365 [56] and a set of sentences describing image type (photo, drawing), image quality (blurry, noise, JPEG compression), weather conditions, etc.
Concerning the urban scene segmentation task, we also built a GT-based sentence set for the WD2 dataset [52]. It contains annotations for 81 classes, from which we selected 74 classes (removing the ambiguous ones most difficult for VLMs to handle, namely ‘unlabeled’, ‘ego-vehicle’, ‘overlay’, ‘out-of-roi’, ‘static’, ‘dynamic’, ‘curb terrain’). The sentences have the form {“An image showing a <class>”}, where <class> was replaced each time by one of the 74 classes selected. While this allows us to generate GT relevance scores for each image in WD2, this set limits our methods in selecting only from failure causes related to the presence of these classes in the image and hence they do not cover all the possible failure reasons that may explain the model’s behavior.
LLM-generated. We also consider a set of sentences generated by a large language model (LLM). We provided GPT-3.5 with the context and the sentence structure and prompted it to provide lists of sentences related to weather conditions, vehicles, urban objects, time of day and road conditions. For each of them we used either the sentence form {“An image taken in/at <weather/road condition or time of day>”} or {“An image showing a/an <vehicle/urban object>”}, where the template is filled by the LLM. This yields to a set of 1016 sentences, including most of the ones we design manually in the User-defined set. On the other end, the limitation of this approach is the noise in the set, which can also include sentences hardly helpful for the task at end, e.g.An image taken on an unpredictable/eclectic/dystopian road” or “An image taken in a corn maze/trick-or-treat/pajama weather”.
C.2 Defining the ground-truth ${\mathcal{S}}_{\beta}^{*}$ for LBEE
To evaluate the predicted sentence set ${\mathcal{R}}_{\mathcal{S}}\subset{\mathcal{S}}$ output by an LBEE approach, we must define a set of ground-truth sentences ${\mathcal{S}}^{*}\subset{\mathcal{S}}$ to compare our predictions to—namely, ground-truth error mode descriptions. To derive such set, we consider the function $\omega_{\theta}:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}$ that measures the model’s performance on each image, denoted by $\omega_{\theta}({\bf x}_{i})$. This function can be the accuracy in the case of a classification task, the mIoU (mean Intersection over Union) score for semantic segmentation, or any other task-specific metric. Given the set of images ${\mathcal{X}}$, the average value of this measure over the entire dataset
$\omega^{\text{avg}}_{\theta}=\frac{1}{\lvert{\mathcal{X}}\rvert}\sum_{{\bf x}_{m}\in{\mathcal{X}}}\omega_{\theta}({\bf x}_{m})\enspace,$ (6)
is the overall performance of ${\mathcal{M}}_{\theta}$ on ${\mathcal{X}}$.
[…]
The set of images associated with a sentence $s_{n}$ is defined as
${\mathcal{X}}_{s_{n}}=\{{\bf x}_{m}\in{\mathcal{X}}|\Gamma({\bf x}_{m},s_{n})=1\}\enspace.$ (7)
Then, for each sentence we compute a score by averaging the model performance over the associated images
$\omega_{\theta}^{s_{n}}=\frac{1}{\lvert{\mathcal{X}}_{s_{n}}\rvert}\sum_{{\bf x}_{m}\in{\mathcal{X}}_{s_{n}}}\omega_{\theta}({\bf x}_{m})\enspace.$ (8)
We use this quantity to represent the sentence’s hardness score (according to the measure $\omega_{\theta}$) and we consider a sentence as describing a failure mode if its value is below the global average $\omega^{\text{avg}}_{\theta}$ by at least a margin $\beta$. Therefore, the desired output ${\mathcal{S}}^{*}$ is the collection of all of these sentences:
${\mathcal{S}}^{*}=\{s_{n}\in{\mathcal{S}}|\,\omega_{\theta}^{s_{n}}<\omega_{\theta}^{\text{avg}}-\beta\}\enspace.$ (9)
Varying $\beta$ allows to evaluate the methods with different severity levels for the desired error descriptions, i.e., a higher $\beta$ restricts the desired set ${\mathcal{S}}^{*}$ to more challenging sentences (see Fig. 10).
C.3 Pseudo-GT image-sentence relevance
When we do not have access to GT information, we rely on VQA [1] to determine whether an image $x_{m}\in{\mathcal{X}}$ can be associated with a sentence $s_{n}$ (description) or not. These binary relevance scores $\Gamma(x_{m},s_{n})$ are necessary to define the GT sentence set ${\mathcal{S}}^{*}$ for LBEE (see Sec. C.2) but also to derive the sentence coverage ratio in the cluster (CR defined in Sec. 5).
In particular, we use OFA [47] and LLaVA [27] with their publicly available pre-trained weights—but any other (possibly more recent) model can be used instead. We do not simply use OFA and LLaVA alone, but we also combine them: combining the CLIP-based LLaVA with non CLIP-based OFA representation allow us to significantly decrease the CLIP bias from the pseudo-GT relevance scores.
To proceed with VQA, we first turn each sentence $s_{n}\in{\mathcal{S}}$ into a question $q_{n}$ such that the expected answer is yes or no. For example the sentence "An image taken at night" was turned into the prompt: “Was the image taken at night? Reply simply with ‘yes’ or ‘no’”. Then we provide VQA model with the image-question pairs $(x_{i},q_{n})$ and turn the yes/no answer into a binary score that is assigned to $\Gamma(x_{i},s_{n})$.
Table 4: Comparison of pseudo-ground truth annotations created using LLaVA and OFA with manual ground truth annotations in WD2.
[…]
OFA 0.799 0.489 0.852 0.148 0.511
LLaVA 0.731 0.623 0.696 0.304 0.377
OFA | LLaVA 0.701 0.735 0.642 0.358 0.265
OFA & LLaVA 0.829 0.377 0.905 0.095 0.623
Comparing VQA pseudo-GT to manually annotated GT. Given the key role that VQA plays in our experimental validation, we carry out an experimental analysis to compare the pseudo-GT image-sentence associations obtained with LLaVA [27] and/or OFA [47] to manually annotated GT. To this end, we rely on the GT annotations from WD2 [52] and the derived 74 sentences (described in Sec. C.1) from which we compute the GT scores for $\Gamma(x_{i},s_{n})$. Turning these sentences into questions, we can gather the pseudo-GT relevance scores $\Gamma_{\text{OFA}}(x_{i},s_{n})$ and $\Gamma_{\text{LLAVA}}(x_{i},s_{n})$.
We evaluate the accuracy of these pseudo-GT associations independently, as well as combined with AND/OR logical operations (OFA & LLaVA and OFA | LLaVA, respectively) and report our results in Tab. 4. Combining the output from both models yields the highest accuracy with respect to the ground-truth. Moreover, the accuracy of both models achieves an accuracy above 70% of the GT performance, with LLAVA hallucinating more (FP) and OFA missing more associations (FN). In light of this study, we use OFA & LLaVA to create pseudo-GT sentence-image relevance scores whenever GT information is not available. Note that the overhead of this computation does not affect the methods themselves, as it is used for evaluation only.
Appendix D Design choices
In this section we detail our default parameter setting. We use Open-CLIP [17] trained on LAION-2B [39] as our default visual-textual embedding space, but one could use other VLM models, such as BLIP [24], etc. To generate the GT sentence set ${\mathcal{S}}^{*}_{\beta}$, we use $\beta=o*\omega_{\theta}^{\text{std}}$, where $\omega_{\theta}^{\text{std}}$ is the standard deviation of $\omega_{\theta}$ computed over ${\mathcal{X}}$ with $o=0.2$ as default value. We set the number of hard and easy clusters to $C=15$ for large datasets (when we analyze the full set, namely urban scene segmentation and the unsupervised analyses on the NICO++{++}+ + datasets) and to $C=5$ for small sets, namely when we analyze the dataset for each class independently (per-class analyses of NICO++{++}+ + and ImageNet). We set the number of retained sentence per cluster for all methods to $K=3$.
In the unsupervised cases, we split ${\mathcal{X}}$ according to $t_{\varphi}^{h}=\varphi_{\theta}^{\text{avg}}+a*\varphi_{\theta}^{\text{std}}$ and $t_{\varphi}^{e}=\varphi_{\theta}^{\text{avg}}-a*\varphi_{\theta}^{\text{std}}$, where $\varphi_{\theta}$ is the model’s output entropy and we set $a=0.2$ as default value. In the per-class analysis using NICO++{++}+ + datasets, we use class probabilities $\rho_{\theta}$ and split according to $t_{\rho}^{h}=\rho_{\theta}^{\text{avg}}-a*\rho_{\theta}^{\text{std}}$ and $t_{\rho}^{e}=\rho_{\theta}^{\text{avg}}-a*\rho_{\theta}^{\text{std}}$. For ImageNet, the splitting is done according to the ranking, with the hard set including both false positives (the GT class not ranked first) and false negatives (an image not labeled with the current class where the model ranked the current class as first)—as described in Appendix A.
TopS SetDiff PDiff FPDiff
taken in autumn       “taken in autumn       “sailboat in the water    “sailboat in the water
with waterways   “with waterways       “sailboat in the rocks       “sailboat
plants in the water   “plants in the water   “goose in the water   “birds in the water
TopS   SetDiff   PDiff   FPDiff
flowers       “plants       “flowers in the rocks       “flowers in the rocks
plants   “flowers in the rocks       “plants in the rocks       “plants in the rocks
flowers in the rocks   “flowers in the water   “monkey in the water   “flowers in the water
TopS   SetDiff   PDiff   FPDiff
train       “train       “ train in the rocks       “ train in the rocks
train in the rocks   “train in the rocks       “bus in the rocks       “bus in the rocks
bus in the rocks   “bus in the rocks   “train in the grass   “train in the grass
TopS   SetDiff   PDiff       FPDiff
cactus       “cactus in the rocks       “cactus in the rocks       “cactus in the rocks
cactus in the rocks   “cactus in the water       “cactus in the water       “cactus in the water
cactus in the grass       “plants in the rocks   “cactus   “cactus
TopS   SetDiff   PDiff   FPDiff
truck taken at sunset       “truck taken at sunset       “truck taken at sunset       “truck taken at sunset
bus taken at sunset   “bus taken at sunset       “bus taken at sunset       “bus taken at sunset
train taken at sunset   “train taken at sunset   “truck in the water   “truck in the water
Figure 12: Five out of 15 hard clusters from NICO++{++}+ + 85 (unsupervised case) explained by different methods.
Appendix E Further qualitative analyses
First, we provide in Tab. 5 an extended version of Tab. 1 Then we show qualitative examples for NICO++{++}+ + 85 in Fig. 12, for ImageNet 1K in Figs. 13 and 14 and for WD2 in Fig. 15,
[…]
WD2 IDD ACDC
Sentences (An image …) ${\mathcal{S}}^{*}$ TS PD SD FP ${\mathcal{S}}^{*}$ TS PD SD FP ${\mathcal{S}}^{*}$ TS PD SD FP
“taken at night” + + + + + + + +
“taken in the evening” + + + + + +
“taken at dusk” + + + + + +
“taken at dawn” $\times$ + + + $\times$
“taken in a stormy weather” + + + $\times$ $\times$ +
“taken in a foggy weather” + $\times$ + + + +
“taken in a rainy weather” + + + $\times$ + + +
“taken in a snowy weather” + + + $\times$ $\times$ +
“taken in a windy weather” + + + + $\times$ $\times$ + +
“taken in a dull weather” + + + + $\times$ $\times$ + + + +
“taken in winter” + + $\times$ $\times$ +
“taken in autumn” $\times$ + $\times$
“with reflection on the road” + + + + + +
“with shadows on the road” $\times$ + $\times$ + +
“with water on the road” + $\times$ $\times$ + +
“with mud on the road” $\times$ + + $\times$ + + $\times$
“with branches on the road” + + $\times$ + + +
“with traffic cone on the road” + + +
“with obstacle on the road” + + $\times$ + $\times$
“with road barrier on the road” $\times$ + $\times$ + + + + + +
“with rail track on the road” $\times$ + + + + + +
“with rocks on the road” $\times$ + + $\times$ $\times$
“showing a highway scene” $\times$ + $\times$ + + $\times$ + + + +
“showing an industrial scene” $\times$
“showing construction site” $\times$ + + + $\times$ + +
“showing sub-urban scene” $\times$ + + $\times$ + +
“showing agricultural field” $\times$ $\times$
“showing a round-about” $\times$ $\times$
“with rickshaw on the road” $\times$ + + + + $\times$ + + + +
“with motorbike on the road” $\times$ + + $\times$ + +
“with bike on the road” $\times$ $\times$ + + $\times$ +
“with vehicle on the road” $\times$ + $\times$ + $\times$ + + + +
“with bus on the road” $\times$ $\times$ + + + $\times$ +
“with tram on the road” $\times$ $\times$ + $\times$ + + +
“with jeep on the road” $\times$ + + + $\times$ $\times$
“with animal on the road” $\times$ + + + + $\times$
“with bird on the road” $\times$ $\times$
“with people on the road” $\times$ + + $\times$ + + $\times$ +
“with crowded background” $\times$ + $\times$
“with crowded foreground” $\times$ $\times$ + + + $\times$
“with motion blur” + + + + + +
“with overexposure” $\times$ $\times$
“with underexposure” $\times$ + + $\times$ +
“with low contrast” $\times$ + + $\times$ +
“of a tunnel” + + + + + + + + +
“of fences” $\times$ $\times$ + +
“of guard-rail” $\times$ $\times$ $\times$ + + +
“of a traffic jam” + + + $\times$
“of a traffic” $\times$ + $\times$ + + + $\times$ +
“of traffic lights” $\times$ +
“of buildings” $\times$ + $\times$ + $\times$ +
“of a parking” $\times$ + + + + $\times$ +
“of a mountain” $\times$ + + $\times$ $\times$
TopS SetDiff PDiff FPDiff
dome       “taken in a dim weather       “taken in a snowy weather       “taken in a dim weather
church       “taken at dusk       “taken in roof garden       “taken in a foggy weather
taken in tower       “taken at night   “taken in a dim weather   “taken in a rainy weather
TopS SetDiff PDiff FPDiff
taken in pet shop       “taken in pet shop       “Shetland sheepdog       “taken in pet shop
taken in kennel outdoor       “veterinarians office       “taken in pet shop       “veterinarians office
golden retriever       “a thumbnail image   “taken in hospital room   “a chow
TopS SetDiff PDiff FPDiff
typewriter keyboard       “with low contrast       “taken in clean room       “rule
computer keyboard       “rule       “with studio lighting       “hardware store
space bar       “taken in conference room   “taken in entrance hall   “with low contrast
TopS SetDiff PDiff FPDiff
oxcart       “taken in medina       “oxcar       “oxcar
taken in a village       “taken in barndoor       “taken in slum       “taken in medina
taken in farm       “taken in hayfield       “taken in medina       “taken in a rice paddy
TopS SetDiff PDiff FPDiff
taken in bookstore       “taken in archive       “taken in slum       “taken in archive
taken in library indoor       “taken in office       “taken in street       “taken in office
bookshop       “taken in shed   “taken in porch   “taken in ticket booth
Figure 13: Hard clusters from ImageNet 1K validation set tested with the pre-trained Vit-B16 model. Here we show examples containing mainly false negative (not recognized) images from the correct class. From top to bottom we have the classes “church”, “golden retriever”, “computer keyboard”, “ox” and “bookshop”.
TopS SetDiff PDiff FPDiff
kit fox       “mongoose       “mongoose       “mongoose
grey fox       “taken in yard       “a grey fox       “a grey fox
red fox       “taken in wild field   “desert vegetation   “coyote
TopS SetDiff PDiff FPDiff
rock crab       “thumbnail image       “SD generated image       “taken in fishpond
Dungeness crab       “damaged image       “taken in galley       “Dungeness crab
fiddler crab       “taken in alcove   “taken at night   “taken in lagoon
TopS SetDiff PDiff FPDiff
stove       “taken in archive       “stove       “stove
taken in kitchen       “thumbnail image       “dishwasher       “dishwasher
taken in galley       “taken in laundromat   “SD generated image   “taken in kitchen
TopS SetDiff PDiff FPDiff
zucchini       “taken in a farm       “vegetable garden       “vegetable garden
cucumber       “taken in a market outdoor       “taken in a farm       “taken in a farm
vegetable garden       “taken in yard   “taken in market outdoor   “taken in market outdoor
TopS SetDiff PDiff FPDiff
water snake       “green mamba       “thumbnail image       “thumbnail image
ringneck snake       “thumbnail image       “taken in forest broadleaf       “ringneck snake
vine snake       “taken in arch   “taken in a cloudy weather   “green mamba
Figure 14: Hard clusters from ImageNet 1K validation set tested with the pre-trained Vit-B16 model. Here we show examples containing mainly false positives (images from other classes) images from the correct class. From top to bottom we have the classes “red fox”, “rock crab”, “microwave”, “cucumber” and “water snake”.
TopS SetDiff PDiff FPDiff
rainy weather      “rainy weather      “rainy weather      “rainy weather
vehicle on the road     “stormy weather      “stormy weather      “stormy weather
dull weather     “dust on the road     “dull weather     “dull weather
TopS     SetDiff     PDiff     FPDiff
rickshaw on the road      “sub-urban scene      “market      “rickshaw on the road
vehicle on the road     “dull weather      “restaurants      “taken in the morning
obstacle on the road     “low contrast     “with defocus blur     “taken in the afternoon
TopS     SetDiff     PDiff     FPDiff
vehicle on the road      “tunnel      “tunnel      “shadows on the road
car on the road     “guard-rail      “guard-rail      “light reflection on the road
motion blur     “dirt on the road     “snowy weather     “motion blur
TopS     SetDiff     PDiff      FPDiff
vehicle on the road      “snowy weather      “snowy weather      “snowy weather
clear weather     “rainy weather      “rainy weather      “rainy weather
snowy weather      “stormy weather     “stormy weather     “taken in winter
TopS SetDiff PDiff FPDiff
"vehicle on the road"      "mud on the road"      "mountain "      "mud on the road"
"obstacle on the road"      "taken in a windy weather"      "terrain"      "rocks on the road"
"jeep on the road"      "round-about scene"      "forest "      "jeep on the road"
TopS     SetDiff     PDiff     FPDiff
traffic      “taken at night      “taken at night      “taken at night
showing a sub-urban scene     “taken in the evening      “taken at dusk      “taken in the evening
vehicle on the road     “with underexposure     “taken in the evening     “taken in a rainy weather
TopS SetDiff PDiff FPDiff
"tunnel"      "tunnel"      "tunnel"      "tunnel"
"vehicle on the road"      "taken at night"      "taken at night"      "light reflections on the road"
"car on the road"      "with high brightness"      "light reflections on the road"      "with motion blur"
TopS SetDiff PDiff FPDiff
"light reflections on the road"      "taken at night"      "taken at night"      "taken at night"
"taken at night"      "taken in the evening"      "taken at dusk"      "taken at dusk"
"vehicle on the road"      "taken at dusk"      "light reflections on the road"      "light reflections on the road"
TopS     SetDiff     PDiff     FPDiff
vehicle on the road      “taken in winter      “snowy weather      “taken in winter
clear weather     “snowy weather      “taken in winter      “dull weather
taken at daytime     “tunnel     “mountain     “snowy weather
Figure 15: Nine out of 15 hard clusters from WD2 explained by different methods.